OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 - 2x + 5 @ganeshie8 @Loser66

OpenStudy (loser66):

what is the vertex form?

OpenStudy (anonymous):

(x+k)...?

OpenStudy (loser66):

google

OpenStudy (loser66):

hehehe.. I got it, how about you? don't wait for me. I will guide base on your work

OpenStudy (anonymous):

not coming up @Loser66

OpenStudy (loser66):

what do you mean?

OpenStudy (anonymous):

y = a(x-h)^2+k

OpenStudy (anonymous):

i got y=(x-1)^2+4

OpenStudy (anonymous):

Find the exact value by using a half-angle identity. sin (5pi/12)

OpenStudy (loser66):

how and why, show your work

OpenStudy (anonymous):

can you help me on this one, since i answered that

OpenStudy (loser66):

no, no, the previous question is not correct

OpenStudy (loser66):

let do one by one

OpenStudy (anonymous):

i plugged it in

OpenStudy (anonymous):

oh wait i mixed up the signs

OpenStudy (loser66):

good

OpenStudy (anonymous):

y = (x + 1)2 - 4

OpenStudy (loser66):

yup

OpenStudy (anonymous):

can you help me with the other one please?

OpenStudy (loser66):

what is the half-angle of sin?

OpenStudy (loser66):

google, I tell you the trick only. \(\dfrac{5\pi}{12}=\dfrac{(5/6)*\pi}{2}\)